A coin lies on a plane whose inclined is equal to the coefficient of friction. In the horizontal direction along the plane, the coin was imparted with the velocity u. Find the steady-state speed of the coin.


1
Expert's answer
2021-05-07T09:47:15-0400


According to question,

"\theta=tan^{-}\mu"

Initial velocity imparted "=u"

Acceleration in horizontal direction, "a=gsin\theta=gsin(tan^{-}\mu)"

Let the length of incline "=l"

Let steady state speed of coin "=v"

Let time taken to reach steady state speed "=t"

"v=u+at"

"v=u+gsin(tan^{-}\mu)t"